# Optimizing Cycle Life Prediction of Lithium-ion Batteries via a Physics-Informed Model

### v1

@@ 5 Conclusion and Future Directions
Our research presents a novel technique of understanding and predicting battery capacity loss curves using a physics-informed model. Our focus on reconstructing these curves rather than on just cycle life prediction offers distinct advantages by providing more robust and interpretable predictions without sacrificing the accuracy achieved in prior work [12]. Parameterizing the ground truth capacity loss curves by fitting an equation inspired from an Arrhenius Law, we train a self-attention model that recovers these parameters and reconstructs the full capacity curves, achieving similar errors to [12]. This approach is flexible to the definition of end of life, offering the advantage of predicting the cycle at which any percent of capacity is lost.
However, there are a few other directions that potentially further improve upon our work. Firstly, we note that battery cycle data is a time series, and hence utilizing entire time series as prediction inputs is one way to incorporate more information and features. This method could potentially improve results given the correct training policy, hyperparameters, and machine learning architectures.
@@
Acknowledgements
The authors are grateful to Tan Nguyen and Alex Pham for their mentorship and continued guidance. They also thank Lingyun Ding for fruitful discussions and his continued input. This research was conducted as part of the RIPS 2023 REU and is supported by the Institute of Pure and Applied Mathematics. Lastly, a special thanks to Susana Serna for her continued guidance and feedback on navigating this project.
@@ Acknowledgements
Disclosures
C.D Nicolae, S. Sameer, N. Sun, and K. Yan are listed as inventors on a patent application related to this work.

### v2

@@ 5 Conclusion and Future Directions
Our research presents a novel technique of understanding and predicting battery capacity loss curves using a physics-informed model. Our focus on reconstructing these curves rather than on just cycle life prediction offers distinct advantages by providing more robust and interpretable predictions without sacrificing the accuracy achieved in prior work [13]. Parameterizing the ground truth capacity loss curves by fitting an equation inspired from an Arrhenius Law, we train a self-attention model that recovers these parameters and reconstructs the full capacity curves, achieving similar errors to Severson et al. [13]. This approach is flexible to the definition of end of life, offering the advantage of predicting the cycle at which any percent of capacity is lost.
However, there are a few other directions that potentially further improve upon our work. Notably, the battery cycle data we used is a time series, and hence utilizing entire time series as prediction inputs is one way to incorporate more information and features. This method could potentially improve results given the correct training policy, hyperparameters, and machine learning architectures.
@@
Appendix A Feature correlation scores with parameters $\hat{A}$ and $\hat{B}$
This section presents the correlation scores between the candidate features and equation parameters $\hat{A}$ and $\hat{B}$. The scores of the five features with the highest correlation are bolded. A detailed discussion of feature extraction methodology and their physical significance is provided in Section 2.
Feature Name Feature Description $\hat{A}$ $\hat{B}$
DeltaQ_logVar Variance of logarithmic changes in capacity difference 0.46 -0.23
DeltaQ_logMin Minimum logarithmic change in capacity difference 0.48 -0.25
DeltaQ_logMean Mean logarithmic change in capacity difference 0.48 -0.25
DeltaQ_logSkew Skewness of logarithmic changes in capacity difference 0.28 -0.18
DeltaQ_logKurt Kurtosis of logarithmic changes in capacity difference 0.34 -0.20
Slope_capacity_fade_2-100 Slope of capacity fade between cycles 2 and 100 -0.50 0.56
Slope_capacity_fade_91-100 Slope of capacity fade between cycles 91 and 100 -0.57 0.53
QD_Max_2 Maximum discharged capacity relative to initial capacity 0.32 -0.15
QD_2 Discharged capacity measured at the second cycle 0.22 -0.18
Intercept_capacity_fade_2-100 Intercept of the capacity fade between cycles 91 and 100 -0.30 0.26
Intercept_capacity_fade_91-100 Intercept of the capacity fade between cycles 91 and 100 0.15 -0.12
Min_IR Minimum internal resistance measured during cycling -0.20 -0.21
IR_100-2 Difference in internal resistance between cycles 100 and 2 0.19 -0.17
Table 1: Feature correlation scores with $\hat{A}$ and $\hat{B}$
Appendix B Training curves in two-stage training procedure
The following are the training curves after each stage of the two-stage training procedure described in Section 3.4. As expected, the curves after Stage 1 provide a rough estimate of the capacity loss curve, and is refined after Stage 2 of training.
(a) Predicted and true cycle lives after Stage 1
(b) Predicted and true cycle lives after Stage 2
Figure 8: Predicted and true cycle lives after each stage of the two-stage training procedure.
Appendix C Training curves in two-stage training procedure
(a) Predicted versus true capacity curves after stage 1 of training.
(b) Predicted versus true capacity curves after stage 2 of training.
Figure 9: Predicted versus true capacity curves after each stage of the two-stage training procedure.